# Regularizing Effect for a Nonlocal Maxwell-Schrödinger System

### v1

@@ 3 Estimates and Convergence
We are in position to address our crucial a priori estimates, which will provide the basis for our main results.
@@
$\sigma=\begin{cases}\frac{2p}{2p-1}\ \ \mbox{if}\ \ p^{*}>m^{\prime},\\ \frac{1}{r+\theta}\ \ \mbox{if}\ \ r+\theta+1>m^{\prime}.\end{cases}$
@@
$\begin{cases}u_{k}\rightharpoonup u\mbox{ weakly in }W^{1,p}_{0}(\Omega)\cap L^{r+\theta+1}(\Omega),\\ v_{k}\rightharpoonup v\mbox{ weakly in }W^{1,p}_{0}(\Omega),\\ u_{k}\to u\mbox{ a.e. in }\Omega\mbox{ and strongly in }L^{q}(\Omega),\forall 1\leqslant q<\max\{p^{*},r+\theta+1\},\\ v_{k}\to v\mbox{ a.e. in }\Omega\mbox{ and strongly in }L^{q}(\Omega),1\leqslant q<p^{*},\ \end{cases}$ (3.2)
@@
At this point, observe that, in order to complete this proof, it is enough to control $\|u_{k}\|_{L^{m^{\prime}}}$, what will be achieved by means of a tailored test function. With this purpose in mind, recalling that $u_{k}\geqslant 0$, and $v_{k}>0$ a.e. in $\Omega$, we consider
@@
Further, we denote $E_{1}=\{x\in\Omega;\ u_{k}<v_{k}+\varepsilon\}$ and $E_{2}=\{x\in\Omega;\ u_{k}\geqslant v_{k}+\varepsilon\}$, it is obvious that $\Omega=E_{1}\cup E_{2}$, and hence
@@
$\leqslant C\|f\|_{L^{m}}\|u_{k}\|_{L^{m^{\prime}}},$
so that $\|u_{k}\|_{L^{r+\theta+1}}\leqslant C\|f\|_{L^{m}}^{\sigma}$, where
$\sigma=\begin{cases}\frac{2p}{2p-1}\ \ \mbox{if}\ \ p^{*}>m^{\prime},\\ \frac{1}{r+\theta}\ \ \mbox{if}\ \ r+\theta+1>m^{\prime}.\end{cases}$
@@
The proof of (3.2) is standard. For instance, by recalling that $u_{k}\geqslant 0$ and $v_{k}>0$ a.e in $\Omega$, it is enough to combine (3.1), the Relich-Kondrachov theorem, Thm 3.1 and 3.3 in [5], p.13 and 15.
The next result handles the passage to the limit in certain delicate terms with low regularity.
@@
($\mbox{H}_{1}$)-($\mbox{H}_{4}$). Consider $0<\theta<\min\{p-1,\frac{p^{2}}{N-p}\}$, $1<p<N$, and $m>\min\{(p^{*})^{\prime},(r+\theta+1)^{\prime}\}$, and $u_{k},v_{k}$ solution of (2.6) given by Proposition 2.2. Then, up to subsequences relabeled the same.
@@
and then, we fix $\omega_{2}\equiv 1$, $s_{2}=\theta+1$ and $t_{2}=+\infty$, and $(iv)$ is satisfied, see p. 5.4. Further, observe that by taking $\varphi=\frac{1}{\lambda}T_{\lambda}(G_{n}(u_{k}))u_{k}$ as a test function in the first equation of (2.6), by noticing that
@@
and once again, under the notation of Lemma 5.4, we set $\omega_{1}=f$, $t_{1}=m$ and $s_{1}=1$ so that $(iii)$ is satisfied.
@@
$\int_{\{|v_{k}|>n\}}|u_{k}|^{r}|v_{k}|^{\theta+1}\leq C\mbox{ and }\ \{|u_{k}|^{r}\}\ \mbox{is uniformly integrable}$
Since $\{u_{k}\}$ and $\{v_{k}\}$ are bounded in $W^{1,p}_{0}(\Omega)$, by Rellich - Kondrachov we have $u_{k},v_{k}\rightarrow u,v$ in $L^{q}(\Omega)$ for $q<p^{*}$, and so that a.e in $\Omega$ , up to subsequences. Thus, as $g$ and $h$ are Carathéodory functions we also have $g(x,u_{k},v_{k}),h(x,u_{k},v_{k})\rightarrow g(x,u,v),h(x,u,v)$ a.e in $\Omega$.
@@
where we have employed hypotheses $\mbox{H}_{1}$ and $\mbox{H}_{2}$. Moreover, recall that by item $(a)$$\{g(.,u_{k},v_{k})u_{k}\}_{\mathbb{N}}$ is uniformly integrable and $\{v_{k}^{\theta+1}\}_{\mathbb{N}}$ is uniformly integrable since $\theta+1<p^{*}$. In this way, $\{g(.,u_{k},v_{k})\}_{\mathbb{N}}$ is also uniformly integrable and hence by the Vitali convergence theorem we prove (b).
@@
It is clear that $h(x,u_{k},v_{k})v_{k}\to h(x,u,v)v$ a.e. in $\Omega.$ Next, observe that, since $g(.,u_{k},v_{k})u_{k}\to g(.,u,v)u$ in $L^{1}(\Omega)$ and by ($\mbox{H}_{2}$) $c_{1}u_{k}^{r}v_{k}^{\theta+1}\leqslant g(x,u_{k},v_{k})$ a.e. in $\Omega$, we already have that
@@
However, by ($\mbox{H}_{3}$), there holds $h(x,u_{k},v_{k})\leqslant d_{2}u_{k}^{r}v_{k}^{\theta+1}$ a.e. in $\Omega$, so that by the Lebesgue Dominated Convergence Theorem, $h(.,u_{k},v_{k})v_{k}\to h(.,u,v)v$ in $L^{1}(\Omega)$.
@@ 4 Existence and Regularizing Effects
$\|u\|_{L^{r+\theta+1}}\leqslant C\|f\|^{\sigma}\ \ \mbox{and}\ \ \|u\|^{2p}_{W^{1,p}_{0}}+\|v\|^{2p}_{W^{1,p}_{0}}+\int_{\Omega}u^{r}v^{\theta+1}\leqslant C\|f\|^{\sigma+1}_{L^{m}}$ (4.1)
where $u\geqslant 0$, $v\geqslant 0$ a.e. in $\Omega$, $q_{1}<\max\{p^{*},r+\theta+1\}$ and $q_{2}<p^{*}$. Moreover, we have where recall that $\sigma=\begin{cases}\frac{2p}{2p-1}\ \ \mbox{if}\ \ p^{*}>m^{\prime},\\ \frac{1}{r+\theta}\ \ \mbox{if}\ \ r+\theta+1>m^{\prime}.\end{cases}$
Further, by Proposition 5.4 we have
@@
It is easy to check this $p>m_{p}^{*}\iff m<(p^{*})^{\prime}$, so that if $(r+\theta+1)^{\prime}<m<(p^{*})^{\prime}$, by Theorem (1.3) the solution $u\in W^{1,p}_{0}(\Omega)\cap L^{r+\theta+1}(\Omega)$, in the case $u$ is Sobolev regularized. In addition, as $m<(p^{*})^{\prime}\iff m_{p}^{**}<p^{*}$, which implies that immersion $L^{p^{*}}(\Omega)\subset L^{m_{p}^{**}}(\Omega)$ is continuous, so if $r+\theta>p^{*}-1$, we obtain $L^{r+\theta+1}(\Omega)\subset L^{p^{*}}(\Omega)$ it is also a continuous immersion, thus $u$ is Lebesgue regularized.
@@
Since $r+\theta>p^{*}-1\iff(p^{*})^{\prime}>(r+\theta+1)^{\prime}$, and so if $m>(r+\theta+1)^{\prime}$, by Theorem (1.3) we get $u\in W^{1,p}_{0}(\Omega)\cap L^{r+\theta+1}(\Omega)$. Indeed, it is clear that
$\frac{N(r+\theta+1)}{N(p-1)+p(r+\theta+1)}>m\iff r+\theta+1>m_{p}^{**}.$
@@
According to Theorem (1.3) we know that $u\in L^{r+\theta+1}(\Omega)$ and $v\in L^{p^{*}}(\Omega)$, in such a way that $u^{r}\in L^{\frac{r+\theta+1}{r}}(\Omega)$ and $v^{\theta}\in L^{\frac{p^{*}}{\theta}}(\Omega)$. Thus, by hypothesis ($\mbox{H}_{4}$) and so interpolation inequality we have
@@
The idea is to use non-smoothness of the data to discard trivial solutions. First we prove that $u\neq 0$. Indeed, by $u=0$, then by ($\mbox{H}_{2}$) we would have $g(.,u,v)\equiv 0$. Thus, since the pair $(u,v)$ satisfies Definition ($P_{F}$), then we would have $\int_{\Omega}f\varphi=0$ for all $\phi\in C^{\infty}_{c}(\Omega)$ and in particular $f\in L^{\infty}(\Omega)$ what contradicts the fact that $f\in L^{m}(\Omega)$ and $m<\frac{N}{p}.$ In this way, $u$ is nontrivial.
@@ 5 Appendix
For the second case, $21<p<$ 2, by (5.1) we have
@@
Thus, by using that $21<p<$ 2, it is enough to notice that by the Hölder inequality for $\frac{2}{p}$ and $\frac{2}{2-p}$, it is straightforward that

### v2

@@ 3 Estimates and Convergence
We are ready to address our crucial a priori estimates, providing the basis for our main results. Here, the main difficulty lies in handling the mixed integral estimates, which are delicate.
@@
$\sigma=\begin{cases}\frac{2p}{2p-1}\ \ \mbox{if}\ m^{\prime}\geqslant r+\theta+1,\\ \frac{1}{r+\theta}\ \ \mbox{if}\ \ m^{\prime}<r+\theta+1.\end{cases}$
@@
$\begin{cases}u_{k}\rightharpoonup u\mbox{ weakly in }W^{1,p}_{0}(\Omega)\cap L^{r+\theta+1}(\Omega),\\ v_{k}\rightharpoonup v\mbox{ weakly in }W^{1,p}_{0}(\Omega),\\ u_{k}\to u\mbox{ a.e. in }\Omega\mbox{ and strongly in }L^{q}(\Omega),1\leqslant q<\max\{p^{*},r+\theta+1\},\\ v_{k}\to v\mbox{ a.e. in }\Omega\mbox{ and strongly in }L^{q}(\Omega),1\leqslant q<p^{*},\ \end{cases}$ (3.2)
@@
At this point, clearly, it is enough to control $\|u_{k}\|_{L^{m^{\prime}}}$, what will be achieved by means of a tailored test function. With this purpose in mind, recalling that $u_{k}\geqslant 0$, and $v_{k}>0$ a.e. in $\Omega$, we denote
@@
Further, by denoting $E_{1}=\{x\in\Omega;\ u_{k}<v_{k}+\varepsilon\}$ and $E_{2}=\{x\in\Omega;\ u_{k}\geqslant v_{k}+\varepsilon\}$, it is obvious that $\Omega=E_{1}\cup E_{2}$, and hence
@@
$\leqslant C\|f\|_{L^{m}}\|u_{k}\|_{L^{m^{\prime}}}.$ (3.7)
Now, notice that if $m^{\prime}\geqslant r+\theta+1$, since $m>\min\{(r+\theta+1)^{\prime},(p^{*})^{\prime}\}$ we have $p^{*}>m^{\prime}$, so that, by (3.4) we have $\|u_{k}\|_{W^{1,p}_{0}}\leqslant C\|f\|_{L^{m}}^{\frac{1}{2p-1}}$. Then, by (3.7), we get
$\|u_{k}\|_{L^{r+\theta+1}}\leqslant C\|f\|_{L^{m}}^{\sigma},\mbox{ where }\sigma=\begin{cases}\frac{2p}{2p-1}\ \ \mbox{if}\ \ m^{\prime}\geqslant r+\theta+1,\\ \frac{1}{r+\theta}\ \ \mbox{if}\ \ m^{\prime}<r+\theta+1.\end{cases}$
@@
The proof of (3.2) is standard. For instance, by recalling that $u_{k}\geqslant 0$ and $v_{k}>0$ a.e in $\Omega$, it is enough to combine (3.1), the Relich–Kondrachov theorem, Thm. 3.1 and Thm. 3.3 in [5].
In the next result, we address the passage to the limit for the mixed nonlinearities and related terms. To handle the most delicate items, we ground on Lemma 5.4 in the Appendix.
@@
[…]
@@
[…]
@@
[…]
@@
[…]
@@
It is clear that $h(x,u_{k},v_{k})v_{k}\to h(x,u,v)v$ a.e. in $\Omega.$ Next, observe that, since $g(.,u_{k},v_{k})u_{k}\to g(.,u,v)u$ in $L^{1}(\Omega)$ and by ($\mbox{H}_{1}$) $c_{1}u_{k}^{r}v_{k}^{\theta+1}\leqslant g(x,u_{k},v_{k})u_{k}$ a.e. in $\Omega$, we already have that
@@
However, by ($\mbox{H}_{4}$), there holds $h(x,u_{k},v_{k})v_{k}\leqslant d_{2}u_{k}^{r}v_{k}^{\theta+1}$ a.e. in $\Omega$, so that by the Lebesgue Dominated Convergence Theorem, $h(.,u_{k},v_{k})v_{k}\to h(.,u,v)v$ in $L^{1}(\Omega)$.
@@ 4 Existence and Regularizing Effects
$\|u\|_{L^{r+\theta+1}}\leqslant C\|f\|^{\sigma}\ \ \mbox{and}\ \ \|u\|^{2p}_{W^{1,p}_{0}}+\|v\|^{2p}_{W^{1,p}_{0}}+\int_{\Omega}u^{r}v^{\theta+1}\leqslant C\|f\|^{\sigma+1}_{L^{m}},$ (4.1)
where $u\geqslant 0$, $v\geqslant 0$ a.e. in $\Omega$, $q_{1}<\max\{p^{*},r+\theta+1\}$, $q_{2}<p^{*}$, and $\sigma$ is given in Proposition 3.1. Further, by Proposition 5.4 we have
@@
First, observe that as $r+\theta>p^{*}-1$, clearly, $\min\{(r+\theta+1)^{\prime},(p^{*})^{\prime}\}=(r+\theta+1)^{\prime}$.
@@
It is immediate that $p>m_{p}^{*}\iff m<(p^{*})^{\prime}$ and $m<(p^{*})^{\prime}\iff m_{p}^{**}<p^{*}$, see (1.4) p. 1.4. Thence, since $(r+\theta+1)^{\prime}<m<(p^{*})^{\prime}$, and by Theorem 1.3, $u\in W^{1,p}_{0}(\Omega)\cap L^{r+\theta+1}(\Omega)$, in the current case $u$ is Sobolev regularized. Notice that, in particular, since $m^{**}_{p}<p^{*}=\min\{p^{*},r+\theta+1\}$, clearly, $u$ is also Lebesgue regularized.
@@
Now, $(p^{*})^{\prime}\leqslant m<\frac{N(r+\theta+1)}{N(p-1)+p(r+\theta+1)}$. Then, by Theorem 1.3, we already know that $u\in W^{1,p}_{0}(\Omega)\cap L^{r+\theta+1}(\Omega)$ and $u$ cannot be Sobolev regularized. However, it is clear that
$m<\frac{N(r+\theta+1)}{N(p-1)+p(r+\theta+1)}\iff m_{p}^{**}<r+\theta+1.$
@@
According to Theorem (1.3) we know that $u\in L^{r+\theta+1}(\Omega)$ and $v\in L^{p^{*}}(\Omega)$, in such a way that $u^{r}\in L^{\frac{r+\theta+1}{r}}(\Omega)$ and $v^{\theta}\in L^{\frac{p^{*}}{\theta}}(\Omega)$. Thus, by hypothesis ($\mbox{H}_{4}$) and the interpolation inequality we have
@@
[…]
@@ 5 Appendix
For the second case, $1<p<2$, by (5.1) we have
@@
Thus, by using that $1<p<2$, it is enough to notice that by the Hölder inequality for $\frac{2}{p}$ and $\frac{2}{2-p}$, it is straightforward that
@@
Acknowlegments
The first author was partially supported by FEMAT/Brazil grant 01/2023 and Programa de Educação Tutorial/FNDE/CAPES/Brazil grant 0331131595. The second author was partially supported by CAPES/Brazil grant 88887.480897/2020-00, CNPq/Brazil grant 142279/2020-0 and FAPEMIG/Brazil grant APQ-04528-22.